Peterson Papers makes personalized stationery of the highest quality. The company maintains a stock of blank note cards, calling cards, stationery, and envelopes. Customers order online, indicating the product type, personalization (monogram, name), font style, and color. The following schedule is typical of an order of 100 calling cards:

Activity Minutes
Process order 5
Wait for production to begin 60
Pull calling cards from inventory 10
Set up machine for font style and color 5
Print calling cards 30
Inspect cards 4
Wait for packaging 9
Package cards for shipping 2
Wait for pickup by FedEx 60


Calculate the manufacturing cycle efficiency. (Round answer to 1 decimal place, e.g. 5.1%.)

Manufacturing Cycle Efficiency

Entry field with incorrect answer now contains modified data

%
Answer #1

Value added time = 30 min

Throughput time = Process time + inspection time + Move time + Queue time

= 30+4+9+2+60 = 105 min

Manufacturing cycle efficiency = Value added time / Throughput time = 30 / 105 = 28.6%